OpenStudy (anonymous):

Solve (x/3) + 2 > 8. x > 30 x > 18 x > 26 x > 22

OpenStudy (michele_laino):

please add -2 to both sides of your inequality

OpenStudy (anonymous):

X > 18

OpenStudy (michele_laino):

namely, calculate this: \[\frac{ x }{ 3 }+2-2>8-2\] simplify please!

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

good job @Michele_Laino for the explanation

OpenStudy (michele_laino):

@abswag13 thank you!

OpenStudy (anonymous):

Thanks everyone! I appreciate it! (:

OpenStudy (michele_laino):

@Rawr_My_Love143 thank you!

OpenStudy (anonymous):

np @Michele_Laino

OpenStudy (michele_laino):

@Rawr_My_Love143 after the calculus above, you got: \[\frac{ x }{ 2 }>6\] do you agree?

OpenStudy (michele_laino):

oops sorry \[\frac{ x }{ 3 }>6\] @Rawr_My_Love143 do you agree?

OpenStudy (michele_laino):

so, if we multiply the above inequality by 3, both sides, we will get: \[\frac{ x }{ 3 }*3>6*3\] do you agree?

OpenStudy (michele_laino):

@Rawr_My_Love143

OpenStudy (michele_laino):

then your answer: